OpenStudy (anonymous):

I am trying to find the best way to Maximize my "companies" profit. I have two products that I sell. Produce at least 30 of product #1. Produce at least 20 of product #2. Product #1 costs $3 per unit to make. Product #2 $5 per unit to make. The total production cost cannot exceed $340. If Product #1 earns a profit of $10 per unit and Product #2 earns a profit of $15 per unit, find the combination of Product #1 and Product #2 that will maximize profit. Ive written out some inequalities to represent this. X is greater than or equal to 30. And y is greater than or equal to 20.

OpenStudy (anonymous):

Cost of production cannot exceed $340, so, 3x+5y is less than or equal to 340. What combination of X and Y would maximize my profit?

OpenStudy (anonymous):

I really need help on this. Im completely stumped.

OpenStudy (anonymous):

There should be 3 points of intersection, I must graph and shade.

OpenStudy (texaschic101):

3x + 5y <= 340 x <= 30 y <= 20 product 1 (y) costs $3 to make with a $10 profit, leaving you with $7 per unit sold. product 2 (x) costs $5 to make with a $15 profit, leaving you with $10 per unit sold You make more selling product 2. 3x + 5y <= 340 3(30) + 5y <= 340 90 + 5y <= 340 5y <= 340 - 90 5y <= 250 y <= 25 I figure if you make more from selling product 2, then you want to sell the minimum of product 1, which is 30. I cannot tell you the correct way to do this because I do not know. But this way, your total profit is : product 1 (x) .....selling 30 will make you (30 x 7) = 210 product 2 (y)....selling 25 will make you (25 x 10) = 250 250 + 210 = 460 (this is the total after the cost of materials has been subtracted)

OpenStudy (texaschic101):

sorry that I am not much help with the graphing

OpenStudy (texaschic101):

oops...the inequalities at the top should be y >= 30 and x >= 20.

OpenStudy (texaschic101):

@Hero ...are you good at these ??

OpenStudy (texaschic101):

oops....also product 1 is x and product 2 is y

OpenStudy (texaschic101):

@ganeshie8 ...help...am I way off on this ?

OpenStudy (texaschic101):

@satellite73 ....please help

OpenStudy (texaschic101):

I am trying to get someone to come and look at this

OpenStudy (anonymous):

Thanks for the help Texaschic, i really need it. Id appreciate it if anyone was able and willing to answer this.

OpenStudy (texaschic101):

I do believe that (30,25) is one point

OpenStudy (texaschic101):

@Opcode ...help

OpenStudy (texaschic101):

@oldrin.bataku ...help please

OpenStudy (texaschic101):

bump the question

OpenStudy (texaschic101):

@Dhalsim ...bump the question

OpenStudy (texaschic101):

Opcode....did I mess up big time on this one

OpenStudy (texaschic101):

@cwrw238 ...help

OpenStudy (texaschic101):

@Dhalsim ...can you bump the question

OpenStudy (texaschic101):

I am sorry Dhalsim....I tried

OpenStudy (anonymous):

I cant bump for aqnother 35 minutes. And thank you. ill figure it out i hope

hero (hero):

@texaschic101 the profit is $15 per unit for the first product and $10 per unit for the second product. The profit is the result of selling price minus cost. Subtracting cost from profit doesn't really mean anything.

hero (hero):

Also x >=30 y >=20

OpenStudy (texaschic101):

I just thought taking the profit and subtracting the cost of making the item would result in how much you actually make

OpenStudy (texaschic101):

I caught my messing up of the inequality signs...I listed my mistake a few posts above

hero (hero):

Cost was already subtracted from the sales price to produce profit. You can't subtract it again.

OpenStudy (texaschic101):

oh.....I see but you would still make more from selling product 2 wouldn't you ?

OpenStudy (texaschic101):

so wouldn't you want to sell the least amount of product 1 and more of product 2

hero (hero):

There are more test points involved than you think. You have to test all of them to be certain.

OpenStudy (texaschic101):

Is there a certain way to solve this ?

OpenStudy (texaschic101):

Because he is talking about graphing and shading

hero (hero):

Yes, that's what it involves...figuring out the solution region.

OpenStudy (texaschic101):

I am so not good at that :(

hero (hero):

It's not really that difficult.

hero (hero):

We already know x >=30 y >=20 3x + 5y <=340 Graphing that creates the region.

hero (hero):

Graphing that should generate at least three possible test points

OpenStudy (texaschic101):

so is the slope -3/5 ?

hero (hero):

That appears to be correct.

hero (hero):

And the those two points are tested in the profit equation P(x,y) = 10x + 15y

hero (hero):

(80,20) appears to be the correct one.

OpenStudy (texaschic101):

oh...I was getting my numbers out of the cost equation

hero (hero):

I suppose you didn't understand how those points were generated to begin with.

OpenStudy (texaschic101):

3x + 5y <= 340 3(80) + 5(20) <= 340 240 + 100 <= 340 yep...that one works 3x + 5y <= 340 3(30) + 5(50) <= 340 90 + 250 <= 340 yep..that works also How did you get so smart.....I wish there was a pill that a person could take that would make them smart......why not, there is a pill for everything else

OpenStudy (texaschic101):

your points work in this inequality

hero (hero):

Yeah, you're right, I thought I calculated them wrong.

OpenStudy (texaschic101):

Hero...you need to invent a smart pill

hero (hero):

Nah, I think the movie limitless demonstrated why such a pill won't ever be created. People will just overdose on it and half the world's population would just get really sick or die.

OpenStudy (texaschic101):

your probably right....they would overdose trying to get too smart...

OpenStudy (texaschic101):

thanks Hero...your the best :)

hero (hero):

The most competitive people would probably die because they would probably think that IQ is directly proportional to how many pills you consume.

OpenStudy (texaschic101):

lol.......they would overdose trying to be geniuses.....losing their common sense when it comes to how many pills one can take without overdosing

OpenStudy (texaschic101):

I am gonna go now hero....thanks again :)